ROTATIONAL CONFORMERS OF GROUP VI (Cr, Mo, and W) METAL BIS(TOLUENE) SANDW ICH COMPLEXES

\underline{JUNG SUP LEE}, SUDESH KUMARI, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Transition metal bis(arene) sandwich complexes may adopt eclipsed or staggered conformations due to the aromatic ring rotations about the metal-arene axis.^{a}^{,}^{b} In this study, the group VI (Cr, Mo, and W) metal bis(toluene) complexes are synthesized in a laser-ablation molecular beam source, and their rotational conformers are identified by pulsedfield-ionization zero-electron-kinetic-energy (ZEKE) spectroscopy. For Cr-bis(toluene), the ZEKE spectrum shows three distinctive vibrationless (0-0) transitions between the ground electronic states of the neutral and ionic complexes at 42739(5), 42745(5), and 42805(5) cm^{−1}, corresponding to ionization energies of 180^{∘}, 60^{∘}/120^{∘}, and 0^{∘} rotamers. In addition, the spectrum exhibits metal-toluene bending (164, 180, 196, and 223 cm^{−1}) and stretching (278 and 291 cm^{−1}) frequencies of these rotamers. The ground electronic states of the 0^{∘} and 180^{∘} rotamers are ^{1}A_{1} (C_{2v}) and ^{1}A_{g} (C_{2h}) in the neutral form and ^{2}A_{1} (C_{2v}) and ^{2}A_{g} (C_{2h}) in the ionized form, respectively. For the 60^{∘} and 120^{∘} rotamers, the ground states of the neutral molecules are ^{1}A (C_{2}), and those of the corresponding ions are ^{2}A (C_{2}). Through the variation of the molecular beam conditions, the eclipsed conformer (0^{∘}) is determined to be more stable than the staggered ones (180^{∘}, 120^{∘}, and 60^{∘}). Similarly, multiple conformers are identified for the Mo and W complexes.

aB.S. Sohnlein, S. Li, and D.S. Yang, J. Chem. Phys. 123, 214306 (2005); B.S. Sohnlein and D.S. Yang, J. Chem. Phys. 124, 134305 (2006)

bS.Y. Ketkov, H.L. Sezle, and F.G.N. Cloke, Angew. Chem. Int. Ed. 46, 7072 (2007) and references therein.